# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

### v1

@@ 3.1 Conservatively Modified Koury-McAllister-Stewart Algorithm
Our first general class of algorithmic approaches based on modified variants of the KMS algorithm, as summarized in Algorithm 1, can be applied with any instances of the general classes of approximate mixed-precision computing methods to solve the linear systems in Step $\mathbf{5}$ and Step $\mathbf{3}$. In particular, we replace the single invocation of a full-precision linear system solver in the original KMS algorithm with an appropriate iterative approximate mixed-precision linear system solver of interest involving a sequence of steps that lead to full-precision accuracy. To reduce the computation costs as much as possible while still realizing full-precision results, we determine a level of reduced precision (such as single-precision, half-precision, bfloat16, and so on) based on the properties of the linear system that ensures convergence under the approximate mixed-precision computing method of interest to yield full-precision accuracy. As a representative example, we shall focus on the iterative refinement method in which case the selection of the proper reduced precision is based on an estimate of the condition number and the norm of the matrix $\mathbf{P}_{ii}$ for each linear system in Step $\mathbf{5}$, $i\in[m]$, and similarly for the linear system of Step $\mathbf{3}$. Specifically, for each linear system with matrix $\mathbf{P}_{ii}$, we determine the machine precision $\delta$ such that $\delta\kappa(\mathbf{P}_{ii})\|\mathbf{P}_{ii}\|$ is sufficiently small to ensure convergence [20], where $\kappa(\mathbf{P}_{ii})$ denotes the condition number of the matrix $\mathbf{P}_{ii}$, $i\in[m]$.
@@
Once the level of reduced precision has been determined in this manner, the iterative refinement method is used to solve each of the linear systems in Step $\mathbf{5}$ and Step $\mathbf{3}$, generically denoted by $\mathbf{A}\mathbf{x}=\mathbf{b}$. At every step $k$ of the iterative refinement method within each outer-loop iteration $t$ of Algorithm 1, we compute the residual $\mathbf{d}_{k}=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}$ and then solve for $\mathbf{y}_{k}$ in $\mathbf{A}\mathbf{y}_{k}=\mathbf{d}_{k}$ using the reduced precision method. The resulting solution $\mathbf{y}_{k}$ is added to the previous estimate $\mathbf{x}_{k}$ to obtain $\mathbf{x}_{k+1}=\mathbf{x}_{k}+\mathbf{y}_{k}$, and the step index $k$ is incremented. This process of inner-loop steps is repeated until the residual falls below the error of double-precision arithmetic, which is guaranteed by our selection of the reduced precision level [20]. An analogous approach can be followed for any of the approximate mixed-precision computing method alternatives to iterative refinement.
Algorithm 1 Conservatively Modified KMS Algorithm
@@
Input: Irreducible stochastic matrix $\mathbf{P}$
@@
 Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
 Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
 Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using           approximate mixed-precision computing method of interest
 Compute Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}$
 Compute solutions of linear systems $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ for $i=m,\ldots,1$           using approximate mixed-precision computing method of interest, and Obtain $\bm{\pi}^{(t)}$
 Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
@@
3.2 Aggressively Modified Koury-McAllister-Stewart Algorithm
Building on the above conservatively modified KMS algorithm to address the performance bottlenecks of the original KMS algorithm, our second general class of algorithmic approaches based on modified variants of the KMS algorithm, as summarized in Algorithm 2, consists of taking a more aggressive approach in exploiting approximate mixed-precision computing methods to replace the standard linear solvers in Step $\mathbf{5}$ and Step $\mathbf{3}$. In particular, Step $\mathbf{5}$ and Step $\mathbf{3}$ can be computed with fewer iterations at the expense of not realizing full-precision results which is then addressed through appropriated levels of preconditioning and iteration for further adjustments. We shall solely focus on Step $\mathbf{5}$ in this section with the understanding that the same approximate mixed-precision computing approach can be similarly applied in Step $\mathbf{3}$;. alternatively, Step $\mathbf{3}$ can be computed by employing our approximate mixed-precision computing approach in Algorithm 1.
Define the block matrices $\mathbf{D}_{ii}=\mathbf{I}_{n_{i}}-\mathbf{P}_{ii}$, $i\in[m]$; the strictly block-lower-triangular matrices $\mathbf{L}_{ij}=\mathbf{P}_{ij}$, $i>j$, and $\mathbf{L}_{ij}=\mathbf{0}$, $i\leq j$, $i,j\in[m]$; and the strictly block-upper-triangular matrices $\mathbf{U}_{ij}=\mathbf{P}_{ij}$, $i<j$, and $\mathbf{U}_{ij}=\mathbf{0}$, $i\geq j$, $i,j\in[m]$. Further define the block-diagonal matrix $\mathbf{D}=\operatorname*{diag}(\mathbf{D}_{11},\ldots,\mathbf{D}_{mm})$. We then can decompose the matrix $\mathbf{I}-\mathbf{P}$ as
@@ 3.2 Aggressively Modified Koury-McAllister-Stewart Algorithm
Now, define the diagonal matrix $\mathbf{D}^{(t-1)}$ of iteration $t-1$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$, $i\in[m]$. From Step $\mathbf{4}$, we then can write $\mathbf{z}^{(t)}=\bm{\pi}^{(t-1)}\mathbf{D}^{(t-1)}$ which, in combination with the systems of linear equations in Step $\mathbf{5}$ of the original KMS algorithm, leads to $\bm{\pi}^{(t)}=\mathbf{z}^{(t)}\mathbf{L}(\mathbf{D}-\mathbf{U})^{-1}$. Upon substituting the expression for $\mathbf{z}^{(t)}$, we finally have
@@
With this as a starting point, our second general class of modified variants of the KMS algorithm is based on exploiting a combination of (13) together with levels of mixed-precision preconditioning and iteration for Step $\mathbf{5}$. In particular, transposing both sides of (13) yields
@@
from which it is apparent that $\bm{\pi}^{(t)\top}$ can be recast as the solution of a sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its right-hand side. Then the main idea of our aggressive mixed-precision computing method for Step $\mathbf{5}$ is to approximately solve the linear systems in (14) via a fixed number of steps of the mixed-precision Richardson iteration method [13]. In essence, the preconditioned Richardson iteration updates the $k_{t}$-th approximation of $\bm{\pi}^{(t)\top}$ via the fixed-point iteration
@@
where $k_{t}$ denotes the number of mixed-precision Richardson iteration steps in the $t$-th iteration of Algorithm 2 and the matrix $\mathbf{M}^{-1}$ denotes the preconditioner of the iterative method. For example, if $\mathbf{M}\equiv\mathbf{D}^{\top}$, i.e., block-Jacobi preconditioning, the iteration matrix in (15) becomes equal to $\mathbf{D}^{-\top}\mathbf{U}^{\top}$. Throughout the rest of this paper we consider $\mathbf{M}$ to be the product of the LU factors of $\mathbf{D}^{\top}$ obtained using reduced precision. [^1] While not explored in this paper, it is worth mentioning that (15) is equivalent to a linear system with the same iteration matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ and multiple right-hand sides, and thus even further improvements might be possible within the context of our approach; see, e.g., [16, 17].
Algorithm 2 Aggressively Modified KMS Algorithm
Input: Irreducible stochastic matrix $\mathbf{P}$
@@
 Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
 Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
 Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using           approximate mixed-precision computing method of interest
 Compute $\mathbf{D}^{(t-1)}$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=$           $(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$, $i\in[m]$.
 Compute solution of sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and          $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its right-hand side using approximate mixed-precision computing           method of Richardson iteration (15), and Obtain $\bm{\pi}^{(t)}$
 Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
@@ 4 Mathematical Analysis
We now turn to derive a mathematical analysis of the general classes of computational algorithmic approaches of the previous section. Our derivation of an error and convergence analysis is presented first for our conservatively modified KMS algorithmic approach and then for our aggressively modified KMS algorithmic approach, establishing the guaranteed convergence of both of our approaches. We conclude this section with a performance analysis of certain aspects of the computational improvements provided by our exploitation of multi-precision arithmetic in both approaches. The conditions and properties expressed in both Section 2 and the introduction are assumed to hold throughout our analysis.
4.1 Error and Convergence Analysis of Conservatively Modified KMS Algorithm
Our main theoretical result is that Algorithm 1 — our conservatively modified KMS algorithm — is guaranteed to converge with an approximation error that decreases by a factor of $O(\epsilon)$ at each iteration. Recall $\bm{\pi}$ to be the exact stationary distribution of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ and $\bm{\pi}^{(t)}$ to be the estimate of the solution $\bm{\pi}$ in iteration $t$. Assuming the estimate $\bm{\pi}^{(t-1)}$ of $\bm{\pi}$ has $O(\varphi)$ accuracy, we first present a lemma that shows the estimate $\mathbf{s}^{(t-1)}$ in iteration $t-1$ of the vector $\mathbf{s}$ to also have $O(\varphi)$ accuracy and further shows the estimate $\mathbf{R}^{(t-1)}$ in iteration $t-1$ of the matrix $\mathbf{R}$ to have $O(\varphi\epsilon)$ accuracy.

### v2

@@ 3.2 Representative Conservative Example of General Algorithmic Framework
As the initial step comprising mixed-precision mode selection, in order to reduce the computation costs as much as possible while still realizing full-precision results, we determine a level of reduced precision (such as single-precision, half-precision, bfloat16, and so on) based on the numerical properties of the linear system that ensures convergence under the IR method to yield full-precision accuracy. For our representative example employing the IR method, the selection of the proper reduced precision is based on an estimate of the condition number and the norm of the matrix $\mathbf{P}_{ii}$ [24, 25] for each linear system in Step $\mathbf{5}$, $i\in[m]$, and similarly for the linear system of Step $\mathbf{3}$. Specifically, for each linear system generically denoted by $\mathbf{A}\mathbf{x}=\mathbf{b}$, we determine the machine precision $\delta$ such that $\delta\kappa(\mathbf{A})\|\mathbf{A}\|$ is sufficiently small to ensure convergence [26], where $\kappa(\mathbf{A})$ denotes the condition number of the matrix $\mathbf{A}$.
@@
Once the level of reduced precision has been determined in this manner, the IR method is used to solve each of the linear systems in Step $\mathbf{5}$ and Step $\mathbf{3}$, again generically denoted by $\mathbf{A}\mathbf{x}=\mathbf{b}$. In particular, at every step $k$ of the IR method within each outer-loop iteration $t$ of Algorithm 1, we compute the residual $\mathbf{d}_{k}=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}$ and then solve for $\mathbf{y}_{k}$ in $\mathbf{A}\mathbf{y}_{k}=\mathbf{d}_{k}$ using the selected reduced-precision method. The resulting solution $\mathbf{y}_{k}$ is added to the previous estimate $\mathbf{x}_{k}$ to obtain $\mathbf{x}_{k+1}=\mathbf{x}_{k}+\mathbf{y}_{k}$, and the step index $k$ is incremented. This process of inner-loop steps is repeated until the residual falls below the error of double-precision arithmetic, which is guaranteed by our selection of the reduced-precision level [26]. We note that an analogous approach can be followed for any of the iterative approximate mixed-precision computing method alternatives to the IR method.
Algorithm 1 Representative Conservative Example of General Algorithmic Framework
@@
Input: Irreducible stochastic matrix $\mathbf{P}$ with structural form (2)
@@
$\mathbf{1.}$ Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
$\mathbf{2.}$ Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
$\mathbf{3.}$ Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using approximate mixed-           precision computing method of interest
$\mathbf{4.}$ Compute Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}$
$\mathbf{5.}$ Compute solutions of linear systems $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ for $i=m,\ldots,1$ using approximate           mixed-precision computing method of interest, and Obtain $\bm{\pi}^{(t)}$
$\mathbf{6.}$ Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
@@
3.3 Representative Aggressive Example of General Algorithmic Framework
The second representative example of a general application of our algorithmic framework within the context of the KMS algorithm also consists of replacing each invocation of a full-precision linear system solver in Step $\mathbf{5}$ and Step $\mathbf{3}$ with a combination of the selection of an appropriate level of mixed-precision supported by the available computer architecture and the selection of any appropriate instance of the general classes of iterative approximate numerical methods that may not necessarily jointly lead to full-precision accuracy. Specifically, as a primary difference with Section 3.2, we take a more aggressive approach in this representative example of our general algorithmic framework by exploiting the general classes of iterative approximate mixed-precision numerical methods such that each occurrence of a linear system can be computed through a sequence of fewer iterations at the expense of not realizing full-precision results, which is then addressed through appropriate levels of preconditioning and iteration for further adjustments. We summarize such a representative aggressive example in Algorithm 2 where the RI method is chosen for the iterative approximate numerical method of our general mathematical framework and is solely applied to Step $\mathbf{5}$, with the understanding that the same approximate mixed-precision numerical method can be applied in Step $\mathbf{3}$; alternatively, Step $\mathbf{3}$ can be computed by employing our approximate mixed-precision computing approach in Algorithm 1.
More formally, define the block matrices $\mathbf{D}_{ii}=\mathbf{I}_{n_{i}}-\mathbf{P}_{ii}$, $i\in[m]$; the strictly block-lower-triangular matrices $\mathbf{L}_{ij}=\mathbf{P}_{ij}$, $i>j$, and $\mathbf{L}_{ij}=\mathbf{0}$, $i\leq j$, $i,j\in[m]$; and the strictly block-upper-triangular matrices $\mathbf{U}_{ij}=\mathbf{P}_{ij}$, $i<j$, and $\mathbf{U}_{ij}=\mathbf{0}$, $i\geq j$, $i,j\in[m]$. Further define the block-diagonal matrix $\mathbf{D}=\operatorname*{diag}(\mathbf{D}_{11},\ldots,\mathbf{D}_{mm})$. We then can decompose the matrix $\mathbf{I}-\mathbf{P}$ as
@@ 3.3 Representative Aggressive Example of General Algorithmic Framework
Now, define the diagonal matrix $\mathbf{D}^{(t-1)}$ of iteration $t-1$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$, $i\in[m]$. From Step $\mathbf{4}$, we then can write $\mathbf{z}^{(t)}=\bm{\pi}^{(t-1)}\mathbf{D}^{(t-1)}$ which, in combination with the systems of linear equations in Step $\mathbf{5}$, leads to $\bm{\pi}^{(t)}=\mathbf{z}^{(t)}\mathbf{L}(\mathbf{D}-\mathbf{U})^{-1}$. Upon substituting the expression for $\mathbf{z}^{(t)}$, we finally have
@@
With this as a starting point, the second representative example of a general application of our algorithmic framework within the context of the KMS algorithm is based on exploiting a combination of (13) together with levels of mixed-precision preconditioning and iteration for Step $\mathbf{5}$. In particular, transposing both sides of (13) yields
@@
from which it is apparent that $\bm{\pi}^{(t)\top}$ can be recast as the solution of a sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its right-hand side. Then the main idea of the representative aggressive example of our general algorithmic framework for Step $\mathbf{5}$ is to approximately solve the linear systems in (14) via a fixed number of steps of the mixed-precision RI method [22]. In essence, the preconditioned RI updates the $k_{t}$-th approximation of $\bm{\pi}^{(t)\top}$ via the fixed-point iteration
@@
where $k_{t}$ denotes the number of mixed-precision RI steps in the $t$-th iteration of Algorithm 2 and the matrix $\mathbf{M}^{-1}$ denotes the preconditioner of the iterative method. For example, if $\mathbf{M}\equiv\mathbf{D}^{\top}$, i.e., block-Jacobi preconditioning, the iteration matrix in (15) becomes equal to $\mathbf{D}^{-\top}\mathbf{U}^{\top}$. Throughout the rest of this paper we consider $\mathbf{M}$ to be the product of the LU factors of $\mathbf{D}^{\top}$ obtained using reduced precision. [^1] While not explored in this paper, it is worth mentioning that (15) is equivalent to a linear system with the same iteration matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ and multiple right-hand sides, and thus even further improvements might be possible within the context of our approach; see, e.g., [27, 28].
Algorithm 2 Representative Aggressive Example of General Algorithmic Framework
Input: Irreducible stochastic matrix $\mathbf{P}$ with structural form (2)
@@
$\mathbf{1.}$ Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
$\mathbf{2.}$ Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
$\mathbf{3.}$ Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using approximate mixed-           precision computing method of interest
$\mathbf{4.}$ Compute $\mathbf{D}^{(t-1)}$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$,           $i\in[m]$
$\mathbf{5.}$ Compute solution of sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its           right-hand side using approximate mixed-precision computing method of RI in (15), and Obtain $\bm{\pi}^{(t)}$
$\mathbf{6.}$ Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
@@ 4 Mathematical Analysis
We now turn to derive a mathematical analysis of the two representative examples of our general algorithmic framework of numerical methods from the previous section. Our derivation of a theoretical error and convergence analysis is presented first for the representative conservative example of our general mathematical framework and then for the representative aggressive example of our general mathematical framework, establishing the guaranteed convergence and the rate of convergence for both representative examples. We conclude this section with a performance analysis of certain aspects of the computational improvements provided by our exploitation of multi-precision arithmetic in both representative examples of our algorithmic framework. The conditions and properties expressed in both Section 2 and the introduction are assumed to hold throughout our mathematical analysis.
4.1 Error and Convergence Analysis of Representative Conservative Example
Our main theoretical result is that Algorithm 1 — a representative conservative example of our general algorithmic framework — is guaranteed to converge with an approximation error that decreases by a factor of $O(\epsilon)$ at each iteration. Recall $\bm{\pi}$ to be the exact stationary distribution of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ and $\bm{\pi}^{(t)}$ to be the estimate of the exact solution $\bm{\pi}$ in iteration $t$. Assuming the estimate $\bm{\pi}^{(t-1)}$ of $\bm{\pi}$ has $O(\varphi)$ accuracy, we first present a lemma that shows the estimate $\mathbf{s}^{(t-1)}$ of the vector $\mathbf{s}$ in iteration $t-1$ to also have $O(\varphi)$ accuracy and further shows the estimate $\mathbf{R}^{(t-1)}$ of the matrix $\mathbf{R}$ in iteration $t-1$ to have $O(\varphi\epsilon)$ accuracy. Recalling $\mathbf{v}_{i}$ to be the left eigenvector of $\mathbf{P}_{ii}$ which corresponds to its dominant eigenvalue $\lambda_{i_{1}}=1-{g}_{i}\epsilon+o(\epsilon)$, $i\in[m]$, the lemma also shows that there is an $O(\epsilon)$ norm difference between the exact stationary distribution $\bm{\pi}$ and the corresponding eigenvectors $\mathbf{v}_{i}$ of $\mathbf{P}_{ii}$ multiplied by an $O(1)$ constant.